OpenStudy (xmistermayhem):

What is the equation of the ellipse with foci (0, 3), (0, -3) and co-vertices (1, 0), (-1, 0)?

OpenStudy (jdoe0001):

any ideas on the center of it?

OpenStudy (xmistermayhem):

I guess 0,0 ?? Not sure...

OpenStudy (jdoe0001):

well, let's see  so... yes, is at the origin, or 0,0 well. what's the distance from the center to either focus? what about the minor axis, what's the length of the minor axis?

OpenStudy (xmistermayhem):

3 to each foci, 1 is the minor axis?

OpenStudy (jdoe0001):

yeap.. so let's see now.. one sec

OpenStudy (xmistermayhem):

Oki doki

OpenStudy (jdoe0001):

is a vertical ellipse, meaing the major axis is over the y-axis, that is "a" is under the fraction with the "y" then so.... one sec

OpenStudy (jdoe0001):

so. the distance from the focus to the center, is "c" and \(\bf \cfrac{(x-{\color{brown}{ h}})^2}{{\color{purple}{ b}}^2}+\cfrac{(y-{\color{blue}{ k}})^2}{{\color{purple}{ a}}^2}=1 \qquad center\ ({\color{brown}{ h}},{\color{blue}{ k}})\qquad vertices\ ({\color{brown}{ h}}, {\color{blue}{ k}}\pm a) \\ \quad \\ \cfrac{(x-{\color{brown}{ 0}})^2}{{\color{purple}{ 1}}^2}+\cfrac{(y-{\color{blue}{ 0}})^2}{{\color{purple}{ a}}^2}=1 \qquad center\ ({\color{brown}{ 0}},{\color{blue}{ 0}}) \\ \quad \\ c=\sqrt{a^2-b^2}\implies 3=\sqrt{a^2-1^2}\impliedby \textit{solve for "a"}\)

OpenStudy (jdoe0001):

so anyhow \(\bf \cfrac{(x-{\color{brown}{ h}})^2}{{\color{purple}{ b}}^2}+\cfrac{(y-{\color{blue}{ k}})^2}{{\color{purple}{ a}}^2}=1 \qquad center\ ({\color{brown}{ h}},{\color{blue}{ k}})\qquad vertices\ ({\color{brown}{ h}}, {\color{blue}{ k}}\pm a) \\ \quad \\ \cfrac{(x-{\color{brown}{ 0}})^2}{{\color{purple}{ 1}}^2}+\cfrac{(y-{\color{blue}{ 0}})^2}{{\color{purple}{ a}}^2}=1 \qquad center\ ({\color{brown}{ 0}},{\color{blue}{ 0}}) \\ \quad \\ c=\sqrt{a^2-b^2}\implies 3=\sqrt{a^2-1^2}\implies 3^2=a^2-1 \\ \quad \\ 9+1=a^2\implies \sqrt{10}=a\qquad thus \\ \quad \\ \cfrac{(x-{\color{brown}{ 0}})^2}{{\color{purple}{ 1}}^2}+\cfrac{(y-{\color{blue}{ 0}})^2}{{\color{purple}{ a}}^2}=1\implies \cfrac{x}{{\color{purple}{ 1}}^2}+\cfrac{y}{{\color{purple}{ (\sqrt{10})}}^2}=1 \\ \quad \\ \cfrac{x}{1}+\cfrac{y}{10}=1\implies x+\cfrac{y}{10}=1\)